Two point charges form a right-angled triangle with the point A at the origin. Their charges are Q2 = 6 x 10-9 C = 6 nC and Q3 = -3 x 10-9 C = -3 nC. The distance between point A and Q2 is 5 x10-2 m and the distance between point A and Q3 is 3 x 10-2 m. What is the net electric field measured at A from the two charges?


1
Expert's answer
2020-07-17T09:21:28-0400

As per the question,

The given charges are "Q_2=6\times 10^{-9}C"

Distance between A and "Q_2" "(r_2)=5\times 10^{-2}m"

"Q_3=-3\times 10^{-9}C"

Distance between A and "Q_3" "(r_3)=3\times 10^{-2}m"

A is at the origin, let "Q_2" is on the x axis and "Q_3" is on the y-axis.

Hence,"E_i=\frac{Q_2}{4\pi \epsilon_o r_2^2}\hat{i}=9\times 10^9\times\frac{6\times 10^{-9}C}{(5\times 10^{-2}m)^2}\hat{(-i)}"

"=\frac{-54}{25\times 10^{-4}}\hat{i}=-2.16\times 10^4\hat{i}"

"E_j=\frac{Q_3}{4\pi \epsilon_o r_3^2}\hat{j}=9\times 10^9\times\frac{-3\times 10^{-9}C}{(3\times 10^{-2}m)^2}\hat{(-j)}"

"=3\times 10^4\hat{j}"

"E_{net}=-2.16\times 10^4\hat{i}+3\times 10^4\hat{j}"

"|E_{net}|=\sqrt{E_i^2+E_j^2} =3.7\times 10^4 N/C"